Roadside Markets has a 6.75 percent coupon bond outstanding that matures in 6 years. The bond pays interest semiannually. What is the market price per bond if the face value is $1,000 and the yield to maturity is 7.70 percent? A. $1,046 B. $955 C. $899 D. $983
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

For this question,

C = 6.75% * $1000/2 = $33.75 (semi-annually),

n = 6 * 2 = 12 semi-annual periods

M = $1000

i = 7.70%/2 = 3.85% (semi-annually)

1- P = 33.75 * - (1+0.0385) 12 0.0385 1000 (1 + 0.0385) 12

P =$319.52 + $635.51

P = $955.03 (Option B)